# Auto-Encoded Supervision for Perceptual Image Super-Resolution

### v1

@@ Appendix E Further discussion on AESOP
Appendix F Understanding AESOP in terms of PD trade-off
Comparison between $\mathcal{L}_{\text{percep}}$ and $\mathcal{L}_{\text{AESOP}}$. The proposed loss $\mathcal{L}_{\text{AESOP}}$ and $\mathcal{L}_{\text{percep}}$ share the characteristic of utilizing a pretrained neural network. However, the purpose of these two losses is fundamentally different. $\mathcal{L}_{\text{percep}}$ falls within the category of perceptual-oriented losses, aiming for improved perception. Thus, it measures the distance between the SR image and the HR image in a space where additional semantics and textural information are expected to reside. This space is obtained by designing a feature extractor (VGG [51]) that has a high-dimensional feature space, that learns to generate additional semantics not present in the raw pixel space.
In contrast, $\mathcal{L}_{\text{AESOP}}$ is fundamentally a reconstruction loss (fidelity loss), but tailored for the perceptual SR task so that it improves fidelity, for a given level of perception. It utilizes an Auto-Encoder which has a low-dimensional bottleneck. The dimension of the bottleneck is identical to the LR image, thus, the Auto-Encoder compresses and removes specific information. Here, the removed information is the perceptual variance, due to the pretraining objective and the architectural choice of the Auto-Encoder. Accordingly, the Auto-Encoded image contains less information than the raw image, contrary to the high-dimensional feature extractor in $\mathcal{L}_{\text{percep}}$. Note that the term Auto-Encoded space of $\mathcal{L}_{\text{AESOP}}$ indicates the space after the decoder, not the bottleneck. Additionally, we clarify that the improvement in perceptual scores by raising $\mathcal{L}_{\text{AESOP}}$ is since it does not hinder the perceptual-oriented guidance provided by perceptual-oriented losses under the SRGAN-framework. $\mathcal{L}_{\text{AESOP}}$ as a standalone loss will not guide towards realism. We keep improvements in perceptual-oriented losses out of the scope of this work.

### v2

@@ Appendix E Further intuition regarding the PD trade-off
Appendix F Additional visualization
Qualitative examples on benchmark datasets. To further illustrate the effectiveness of our method, we present an additional qualitative comparison between AESOP against the baseline method LDL [37] on the bicubic $\times$4 SR task. We provide results of tested methods, AESOP and LDL, on top of the SwinIR backbone (Fig.21 and Fig.22) and the RRDB backbone (Fig.23 and Fig.24). As can be seen, AESOP significantly improves perceptual quality while effectively suppressing visual artifacts observed in the baseline method.
Additional perception-distortion trade-off curves. We provide extensive visualizations of the perception-distortion trade-off curves in Fig.14-20. For CALGAN [48], we present only a single data point rather than the full perception-distortion trade-off curve, as its official weights are not publicly available. Extensive results show that AESOP leads to substantial performance improvements against baselines in terms of the perception-distortion trade-off. Aligning to Tab.9, AESOP also often outperforms CALGAN even without MoE-discriminator proposed in CALGAN. Additionally, we observe that AESOP often results in larger improvements for Swin Transformer-based methods (e.g., SwinIR, DRCT) compared to CNN-based methods (e.g., RRDB). This is likely because these models have greater capacity and benefit more from the enhanced reconstruction guidance provided by AESOP. However, there are instances where AESOP does not always lead to improved performance. Specifically, AESOP often fails to enhance performance on the Manga109 [44] dataset, which is consistent with the unexpected trade-off behaviors observed across most methods in this dataset. This limitation arises because Manga109 consists predominantly of comic images, which typically lack the fine-grained textures found in photorealistic datasets. The absence of such textures poses a challenge for perceptual SR methods, including AESOP, which are specifically designed to enhance and preserve realistic textures. Consequently, without the presence of these detailed textures, AESOP’s advantages in minimizing fidelity bias and preserving perceptual variance are less pronounced, leading to suboptimal performance in this particular dataset.